suppose three test kits are tested at random. let D represent the defective test kit and let N represent the non defective test kit. if we let x be the random variable for the number of defective test kits, construct the probability distribution of the random variable X


1
Expert's answer
2021-04-29T16:58:27-0400

"D" and "N" are equally likely to occur.

Then we have binomial distribution with parameters


"p=\dfrac{1}{2}"


"n=3"


Denote probability for "k" defective test kits as "P(X=k)".

Then


"P(X=k)=\dbinom{n}{k}p^k(1-p)^{n-k}"


Probabilities for each "k" are


"P(X=0)=\dbinom{3}{0}\cdot\left(\dfrac{1}{2}\right)^0\cdot\left(\dfrac{1}{2}\right)^3=" "\dfrac{3!}{0!3!}\cdot\dfrac{1}{8}=\dfrac{1}{8}"


"P(X=1)=\dbinom{3}{1}\cdot\left(\dfrac{1}{2}\right)^1\cdot\left(\dfrac{1}{2}\right)^2=" "\dfrac{3!}{1!2!}\cdot\dfrac{1}{8}=\dfrac{3}{8}"


"P(X=2)=\dbinom{3}{2}\cdot\left(\dfrac{1}{2}\right)^2\cdot\left(\dfrac{1}{2}\right)^1=" "\dfrac{3!}{2!1!}\cdot\dfrac{1}{8}=\dfrac{3}{8}"


"P(X=3)=\dbinom{3}{3}\cdot\left(\dfrac{1}{2}\right)^3\cdot\left(\dfrac{1}{2}\right)^0=" "\dfrac{3!}{3!0!}\cdot\dfrac{1}{8}=\dfrac{1}{8}"


Check the sum of probabilities


"\displaystyle\sum\limits_{k=0}^3P(X=k)=\dfrac{1}{8}+\dfrac{3}{8}+\dfrac{3}{8}+\dfrac{1}{8}=1"


Construct the probability distribution


"\begin{array}{|c|c|c|c|c|}\hline \\X&0&1&2&3\\ \\ \hline \\P(X=k)&\dfrac{1}{8}&\dfrac{3}{8}&\dfrac{3}{8}&\dfrac{1}{8}\\ \\\hline\end{array}"